OpenStudy (anonymous):

Find all numbers for which the rational expression is undefined. (r^3-3r)/(r^2-49)

OpenStudy (anonymous):

\[(r^3-3r)/(r^2-49) = (r)(r^2-3)/(r-7)(r+7) holes=none, VA= + or - 7\]

OpenStudy (anonymous):

to find HA you need to long divide denominator from numerator

OpenStudy (anonymous):

start by factoring out the numerator and denominator. \[r^3-3r=r(r^2-3)/(r-7)(r+7), r \neq \pm7\]

OpenStudy (anonymous):